OpenStudy (jordanloveangel):

What is the value of x? X+12= -7 x=?

OpenStudy (anonymous):

x + 12 = -7 -12 -12 x= -7 - 12 x= -19

OpenStudy (jordanloveangel):

thanx i still have more can u help

OpenStudy (anonymous):

sure what is it

OpenStudy (jordanloveangel):

-15 - X =-6 X =?

OpenStudy (anonymous):

- 15 -x = -6 +15 +15 -x = 9 /-1 /-1 x = -9

OpenStudy (jordanloveangel):

-6 X=42 X=?

OpenStudy (anonymous):

-6x = 42 /-6 /-6 x = -7

OpenStudy (jordanloveangel):

-2 X+8=-16 x=?

OpenStudy (anonymous):

-2x + 8 = -16 -8 -8 -2x = -24 /-2 /-2 x= 12

OpenStudy (jordanloveangel):

3(X- 2) + 5x=18 x=?

OpenStudy (anonymous):

3(x-2) + 5x = 18 3x-6 + 5x = 18 8x - 6 = 18 -6 -6 8x = 12 /8 /8 x = 3/2

OpenStudy (jordanloveangel):

@Hero

OpenStudy (jordanloveangel):

@zepdrix

OpenStudy (jordanloveangel):

@ganeshie8